OpenStudy (anonymous):

Evaluate. \[\lim_{x \rightarrow -\infty} \ln ^{2x}\]

OpenStudy (anonymous):

I don't understand why the answer is neg infinity

OpenStudy (anonymous):

Or did I wrote it wrong, should it be this: \[\lim_{x \rightarrow -\infty}\ln ^{2^{x}}\] ? (This is my class notes, there's a possibility that I wrote the function wrong :/)

zepdrix (zepdrix):

Ahh sorry :D ok game just ended. So what .. what now? Is the 2^x supposed to be inside the log? doesn't really make sense otherwise.

zepdrix (zepdrix):

Do you think it's this one?\[\Large\rm \lim_{x\to-\infty}\ln(2^x)\]or this one?\[\Large\rm \lim_{x\to-\infty}\ln^2(x)\]

zepdrix (zepdrix):

or this? (ln x)^(2x) bahh im confused +_+

OpenStudy (anonymous):

I think is the first one or the third way... Sorry. My notes wasn't clear...

zepdrix (zepdrix):

OK yah the first one makes sense with the answer you provided.

zepdrix (zepdrix):

\[\Large\rm \lim_{x\to-\infty}\ln(2^x)\quad=\quad \ln\left(2^{(-\infty)}\right)\quad *\]So you can think of that 2 as becoming really really really small, because of the negative in the exponent, ya? Remember how Rational was explaining that in a previous problem?

zepdrix (zepdrix):

\[\Large\rm =\ln(0)\quad*\]

zepdrix (zepdrix):

And you recall that the log function is asymptotic at x=0?

zepdrix (zepdrix):

Going down down down, ya?